# Universal Latent Homeomorphic Manifolds: Cross-Domain Representation Learning via Homeomorphism Verification

### v1

@@ A.3 Proof of Theorem 2 (Bi-Lipschitz Sufficiency)
To prove $E$ is a homeomorphism, we must show it is injective, continuous, and has a continuous inverse. We denote the metrics on input and latent spaces as $d_{X}$ and $d_{Z}$.
@@ 1. Injectivity (Prevention of Collapse)
Assume distinct inputs $x,y\in\mathcal{M}_{x}$. Since $d_{X}$ is a metric, $d_{X}(x,y)>0$. Using the lower bound of the hypothesis:
$d_{Z}(E(x),E(y))\geq c_{1}d_{X}(x,y)$ (28)
Since $c_{1}>0$, it follows that $d_{Z}(E(x),E(y))>0\implies E(x)\neq E(y)$. Thus, $E$ is injective.
@@ 2. Continuity of $E$ (Prevention of Tearing)
The upper bound states $d_{Z}(E(x),E(y))\leq c_{2}d_{X}(x,y)$. This is the definition of Lipschitz continuity. For any $\varepsilon>0$, choosing $\delta=\varepsilon/c_{2}$ guarantees that $d_{X}(x,y)<\delta\implies d_{Z}(E(x),E(y))<\varepsilon$.
3. Continuity of $E^{-1}$ (Metric Stability)
Let $u,v\in\mathcal{M}_{z}$ be points in the image. Since $E$ is bijective onto its image, there exist unique $x,y\in\mathcal{M}_{x}$. Rearranging the lower bound:
$c_{1}d_{X}(x,y)$ $\leq d_{Z}(E(x),E(y))$ (29)
$d_{X}(E^{-1}(u),E^{-1}(v))$ $\leq\frac{1}{c_{1}}d_{Z}(u,v)$ (30)
This inequality demonstrates that the inverse function $E^{-1}$ is Lipschitz continuous with constant $L=1/c_{1}$. Since $c_{1}>0$, $L<\infty$, ensuring continuity.
Since $E$ is a continuous bijection with a continuous inverse, it is by definition a homeomorphism. Therefore, enforcing the bi-Lipschitz constraint ($c_{1}>0,c_{2}<\infty$) is a sufficient condition to satisfy the Local Preservation requirement of Theorem 1. ∎
A.4 Proof of Proposition 1 (Metric Verification)
Proof.
We establish that violations of theoretical manifold properties manifest as detectable changes in computable metrics. We demonstrate one direction: how each theoretical failure mode produces a measurable signature.
@@ A.4 Proof of Proposition 1 (Metric Verification)
Part 1: Structural Fragmentation. Manifold unification (Theorem 1) requires the union $\mathcal{Z}_{\text{total}}$ forms a single connected component. The 0-th Betti number $\beta_{0}$ counts connected components via persistent homology. If domains fail to unify, there exist disjoint subsets $\mathcal{A},\mathcal{B}\subset\mathcal{Z}_{\text{total}}$ with no connecting paths in the $\epsilon$-neighborhood graph. This topological disconnection is detected by $\beta_{0}>1$. Note that $\beta_{0}>1$ with small $W_{2}$ may indicate sampling artifacts rather than true fragmentation; thus both metrics must be evaluated jointly.
Part 2: Geometric Misalignment. Let $\mu=\hat{\mathbb{P}}_{z}^{[i]}$ and $\nu=\hat{\mathbb{P}}_{z}^{[j]}$ be latent distributions. If domains occupy disjoint regions with geometric separation $\Delta>0$, meaning $\inf_{z\in\text{supp}(\mu),z^{\prime}\in\text{supp}(\nu)}\|z-z^{\prime}\|\geq\Delta$, then the Wasserstein-2 distance satisfies:
$W_{2}^{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\mathbb{E}_{(z,z^{\prime})\sim\gamma}[\|z-z^{\prime}\|^{2}]\geq\Delta^{2}$
since any coupling $\gamma$ must transport mass across the gap. Thus geometric misalignment produces large $W_{2}$.
@@
Part 3: Manifold Collapse. The bi-Lipschitz lower bound $c_{1}>0$ (Theorem 2) ensures injectivity. If $c_{1}\to 0$, there exist distinct points $x_{q},x_{\text{adv}}$ with $x_{q}\neq x_{\text{adv}}$ but $\|E(x_{q})-E(x_{\text{adv}})\|\to 0$. When these points have different semantic labels ($y_{q}\neq y_{\text{adv}}$), the $\kappa$-nearest neighborhood $\mathcal{N}_{\kappa}(E(x_{q}))$ in latent space will contain semantic outliers. The Trust Score:
$\text{Trust}_{\kappa}(x_{q})=\frac{1}{\kappa}\sum_{z_{j}\in\mathcal{N}_{\kappa}(E(x_{q}))}\mathbbm{1}[y_{j}=y_{q}]$
decreases when semantic outliers enter the neighborhood. Systematic collapse across the manifold produces statistically significant Trust Score reduction below the expected value for an injective mapping.
Part 4: Structural Incoherence. The Continuity metric [41] measures whether $\kappa$-nearest neighbors in the latent space $\mathcal{Z}$ were also neighbors in the observation space $\mathcal{X}$:
$\text{Cont}_{\kappa}=1-\frac{2}{N\kappa(2N-3\kappa-1)}\sum_{i=1}^{N}\sum_{j\in\mathcal{N}_{i}^{\kappa}(\mathcal{Z})}\max(0,r_{\mathcal{X}}(i,j)-\kappa)$
where $r_{\mathcal{X}}(i,j)$ is the rank of $j$ in $i$’s neighborhood in the original space. If the manifold lacks smoothness, with high local variance, spurious neighbors appear in $\mathcal{Z}$ that were distant in $\mathcal{X}$, increasing the penalty terms and reducing Continuity.
@@
Part 5: Inconsistent Mapping. For paired modalities $(x_{i}^{(m)},s_{i}^{(l)})$ representing the same entity, Theorem 1 requires both map to nearby latent coordinates. If this correspondence is violated, the Alignment Error:
$\tau_{a}=\frac{1}{N_{\text{paired}}}\sum_{i=1}^{N_{\text{paired}}}\|E_{x}^{(m)}(x_{i}^{(m)})-E_{s}^{(l)}(s_{i}^{(l)})\|_{2}$
grows proportionally to the extent of misalignment. Large $\tau_{a}$ indicates failure of cross-modal synchronization. ∎
Remark 2.
This proof establishes that theoretical failures produce detectable metric changes. The converse (metric failure implies theoretical violation) holds under regularity conditions and sufficient sampling density, which we validate empirically through the verification protocol.

### v2

@@ A.2 Proof of Theorem 2 (Bi-Lipschitz Sufficiency)
Recall that $E$ is $(c_{1},c_{2})$-bi-Lipschitz if for all $x,y\in\mathcal{M}_{x}$,
$c_{1}d_{X}(x,y)\leq d_{Z}(E(x),E(y))\leq c_{2}d_{X}(x,y).$ (28)
We must show that $E$ is a continuous bijection onto its image with continuous inverse.
@@ 1. Injectivity (Prevention of Collapse)
Let $x,y\in\mathcal{M}_{x}$ be distinct. Since $d_{X}$ is a metric, $d_{X}(x,y)>0$. Using the lower bound in (28):
$d_{Z}(E(x),E(y))\geq c_{1}d_{X}(x,y)>0,$ (29)
where the final inequality holds since $c_{1}>0$. Therefore $E(x)\neq E(y)$, proving $E$ is injective.
@@ 2. Continuity of $E$ (Prevention of Tearing)
The upper bound in (28) states $d_{Z}(E(x),E(y))\leq c_{2}d_{X}(x,y)$, which is precisely the definition of Lipschitz continuity with constant $c_{2}$. For any $\varepsilon>0$, choosing $\delta=\varepsilon/c_{2}$ ensures that
$d_{X}(x,y)<\delta\implies d_{Z}(E(x),E(y))\leq c_{2}d_{X}(x,y)<c_{2}\cdot\frac{\varepsilon}{c_{2}}=\varepsilon.$
Thus $E$ is continuous.
3. Continuity of $E^{-1}$ (Metric Stability).
Let $\mathcal{M}_{z}^{\prime}:=E(\mathcal{M}_{x})\subseteq\mathcal{M}_{z}$, equipped with the subspace metric induced by $d_{Z}$. By definition, $E:\mathcal{M}_{x}\to\mathcal{M}_{z}^{\prime}$ is surjective, and by Paragraph 1 it is injective; hence $E$ is bijective and admits an inverse $E^{-1}:\mathcal{M}_{z}^{\prime}\to\mathcal{M}_{x}$.
Let $u,v\in\mathcal{M}_{z}^{\prime}$. Then there exist unique $x,y\in\mathcal{M}_{x}$ such that $E(x)=u$ and $E(y)=v$. Using the lower bound in (28), we obtain
$c_{1}d_{X}(x,y)$ $\leq d_{Z}(E(x),E(y))=d_{Z}(u,v),$ (30)
[…]
Thus $E^{-1}$ is Lipschitz continuous with constant $1/c_{1}$, and in particular continuous.
Conclusion
Therefore $E:\mathcal{M}_{x}\to E(\mathcal{M}_{x})$ is a homeomorphism onto its image. In particular, applying the same argument to each modality shows that each $E^{[k]}:\mathcal{M}_{x}^{[k]}\to\mathcal{M}_{z}^{[k]}$ is a homeomorphism onto its image. Hence the bi-Lipschitz constraint is sufficient for Condition (i) of Theorem 1. ∎
A.3 Mathematical Justification for Metric Selection
We establish the connection between violations of theoretical manifold properties (Theorems 1 and 2) and detectable signatures in computable metrics. The arguments are one-directional: each theoretical failure mode implies characteristic degradation in at least one metric in our verification pipeline. While these do not constitute formal proofs of equivalence, they provide principled justification for our metric choices.
@@ A.3 Mathematical Justification for Metric Selection
Part 1: Structural Fragmentation. A necessary empirical indicator of successful topological unification is that the sampled latent union
$\mathcal{Z}_{\mathrm{total}}:=\bigcup_{k}\mathcal{M}_{z}^{[k]}$
behaves as a single connected component at an appropriate neighborhood scale. The 0-th Betti number $\beta_{0}$, computed via persistent homology, counts the number of connected components.
If the conditions of Theorem 1 fail such that domains remain topologically disjoint, then for $\epsilon$ in the operating range (or across a stable persistence interval), there exist disjoint subsets $\mathcal{A},\mathcal{B}\subset\mathcal{Z}_{\mathrm{total}}$ with no connecting paths in the $\epsilon$-neighborhood graph. This topological disconnection is detected by $\beta_{0}>1$.
Because finite samples may appear disconnected at very small scales, $\beta_{0}$ is interpreted jointly with geometric metrics rather than in isolation. The combination $\beta_{0}>1$ with high pairwise Wasserstein distances provides strong evidence of genuine structural fragmentation.
Part 2: Geometric Misalignment. Let $\mu=\hat{\mathbb{P}}_{z}^{[i]}$ and $\nu=\hat{\mathbb{P}}_{z}^{[j]}$ be empirical latent distributions for two domains. The Wasserstein-2 distance is defined as
$W_{2}^{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\mathbb{E}_{(z,z^{\prime})\sim\gamma}[\|z-z^{\prime}\|^{2}],$
where $\Pi(\mu,\nu)$ denotes all couplings with marginals $\mu$ and $\nu$.
Suppose the supports are geometrically separated by a margin $\Delta>0$:
$\inf_{z\in\mathrm{supp}(\mu),\;z^{\prime}\in\mathrm{supp}(\nu)}\|z-z^{\prime}\|\geq\Delta.$
Then for any coupling $\gamma\in\Pi(\mu,\nu)$, we have
$\mathbb{E}_{(z,z^{\prime})\sim\gamma}[\|z-z^{\prime}\|^{2}]\geq\Delta^{2},$
since any transport plan must move mass across the separation gap. Therefore,
$W_{2}^{2}(\mu,\nu)\geq\Delta^{2}.$
This shows that when two latent domain supports are separated by a positive margin, the Wasserstein distance must be large. Such geometric separation is a common failure mode of unification—leading to negligible overlap in latent space—and is therefore detectable via $W_{2}$.
@@
Part 3: Manifold Collapse. The bi-Lipschitz lower bound $c_{1}>0$ (Theorem 2) guarantees injectivity. If this bound is violated (i.e., $c_{1}\to 0$), there exist distinct points $x_{q}\neq x_{\mathrm{adv}}$ such that
$\|E(x_{q})-E(x_{\mathrm{adv}})\|\to 0.$
When these points correspond to different semantic labels ($y_{q}\neq y_{\mathrm{adv}}$), the $\kappa$-nearest neighborhood $\mathcal{N}_{\kappa}(E(x_{q}))$ in latent space will contain semantically inconsistent points.
The neighborhood label purity score
$\mathrm{Purity}_{\kappa}(x_{q})=\frac{1}{\kappa}\sum_{z_{j}\in\mathcal{N}_{\kappa}(E(x_{q}))}\mathbbm{1}[y_{j}=y_{q}]$
measures the fraction of neighbors sharing the same label. As label-inconsistent points enter the neighborhood due to collapse, the purity score decreases. Systematic collapse across the manifold yields statistically significant purity degradation relative to a well-separated baseline embedding.
Note that this argument assumes semantic labels correlate with manifold structure, which is reasonable for supervised settings but may not hold universally.
Part 4: Structural Incoherence. The Continuity metric [52] evaluates whether $\kappa$-nearest neighbors in the latent space $\mathcal{Z}$ were also neighbors in the observation space $\mathcal{X}$:
$\mathrm{Cont}_{\kappa}=1-\frac{2}{N\kappa(2N-3\kappa-1)}\sum_{i=1}^{N}\sum_{j\in\mathcal{N}_{i}^{\kappa}(\mathcal{Z})}\max(0,r_{\mathcal{X}}(i,j)-\kappa),$
where $r_{\mathcal{X}}(i,j)$ denotes the rank of $j$ in $i$’s neighborhood in the original space.
While Theorem 1 only requires homeomorphisms (which preserve topology but not necessarily distances), practical embeddings should preserve local geometric structure. If the embedding severely distorts local neighborhoods (e.g., via folding, tearing, or strong metric distortion inconsistent with the bi-Lipschitz upper bound $c_{2}$), points that were distant in $\mathcal{X}$ may appear as neighbors in $\mathcal{Z}$. This increases the penalty terms $\max(0,r_{\mathcal{X}}(i,j)-\kappa)$ and reduces Continuity.
@@
Part 5: Inconsistent Mapping. For multimodal datasets where paired modalities $(x_{i}^{(m)},s_{i}^{(l)})$ represent the same underlying entity, successful unification requires consistent latent mappings. The Alignment Error
$\tau_{a}=\frac{1}{N_{\mathrm{paired}}}\sum_{i=1}^{N_{\mathrm{paired}}}\|E_{x}^{(m)}(x_{i}^{(m)})-E_{s}^{(l)}(s_{i}^{(l)})\|_{2}$
quantifies whether paired samples map to nearby latent coordinates.
Large values of $\tau_{a}$ indicate failure of cross-modal synchronization. While this condition is not explicitly part of Theorems 1 or 2, it is necessary for practical unified reasoning across modalities. High alignment error suggests that either the encoders have not learned the correct cross-modal correspondence, or the modalities contain complementary rather than redundant information requiring separate latent subspaces.
Remark 3 (Connection to Transfer Learning).
When all verification metrics indicate successful unification—specifically, when domains $i$ and $j$ satisfy $\beta_{0}\approx 1$, low pairwise $W_{2}$, high neighborhood label purity, high Continuity, and (for paired data) low Alignment Error—we have empirical evidence that the latent manifolds $\mathcal{M}_{z}^{[i]}$ and $\mathcal{M}_{z}^{[j]}$ are indistinguishable at the resolution of our measurements.
In this regime, a neural network trained on $\mathcal{M}_{z}^{[i]}$ for a fixed task encounters no detectable distributional shift at the resolution of our measurements when applied to $\mathcal{M}_{z}^{[j]}$ (up to finite-sample variability). This justifies direct transfer learning: the same network parameters can be applied to both domains without fine-tuning, as both domains are represented within the same geometric and topological structure.
This empirical transfer principle is a practical consequence of successful manifold unification, though it is not a formal mathematical theorem.